OpenStudy (gabylovesyou):

Which of the following is the slope between the two points (1, 5) and (6, 4)? 1/5 -1/5 5 -5

OpenStudy (gabylovesyou):

@AnimalAin @Algebraic! @andreascott @azee53 @bngaure @baldymcgee6 @Coolsector

OpenStudy (anonymous):

Yeah im online btw ;) do you know slope formula?

OpenStudy (gabylovesyou):

no ..

OpenStudy (anonymous):

\[slope = m = \frac{ y1 - y2 }{ x1 - x2 }\]

OpenStudy (anonymous):

given two points (x1,y1) , (x2,y2)

OpenStudy (baldymcgee6):

thats wrong

OpenStudy (anonymous):

learn it answers -1/5

OpenStudy (gabylovesyou):

its D?

OpenStudy (anonymous):

b

OpenStudy (baldymcgee6):

m = (y2-y1)/(x2-x1)

OpenStudy (baldymcgee6):

\[m = \frac{ 4-5 }{ 6-1 }\]